Show that ∞n=1[(−1)^(n−1)].[(x^n)/(n)] converges if |x| < 1


1
Expert's answer
2021-02-24T06:40:07-0500

"\sum _{n=1}^{\infty \:}\left(-1^{n-1}\right)x^n\frac{1}{n}"

"\mathrm{If\:there\:exists\:an\:}N\mathrm{\:so\:that\:for\:all\:}n\ge N,\:\quad a_n\ne 0\mathrm{\:and\:}\lim _{n\to \infty }|\frac{a_{n+1}}{a_n}|=L:"

"\mathrm{If\:}L<1\mathrm{,\:then\:}\sum a_n\mathrm{\:converges}"

"\mathrm{If\:}L>1\mathrm{,\:then\:}\sum a_n\mathrm{\:diverges}"

"\mathrm{If\:}L=1\mathrm{,\:then\:the\:test\:is\:inconclusive}"

"\left|\frac{a_{n+1}}{a_n}\right|=\left|\frac{\left(-1^{\left(n+1\right)-1}\right)x^{\left(n+1\right)}\frac{1}{\left(n+1\right)}}{\left(-1^{n-1}\right)x^n\frac{1}{n}}\right|"

"\lim _{n\to \infty \:}\left(\left|\frac{\left(-1^{\left(n+1\right)-1}\right)x^{\left(n+1\right)}\frac{1}{\left(n+1\right)}}{\left(-1^{n-1}\right)x^n\frac{1}{n}}\right|\right)"

"\frac{\left(-1^{\left(n+1\right)-1}\right)x^{n+1}\frac{1}{n+1}}{\left(-1^{n-1}\right)x^n\frac{1}{n}}"

"\frac{-1^{n+1-1}x^{n+1}\frac{1}{n+1}}{-1^{n-1}x^n\frac{1}{n}}"

"\frac{-\frac{1}{n+1}x^{n+1}}{-1^{n-1}\cdot \frac{1}{n}x^n}"

"\frac{x^{n+1}\frac{1}{n+1}}{1^{n-1}x^n\frac{1}{n}}"

"\frac{\frac{x^{n+1}}{n+1}}{1^{n-1}\cdot \frac{1}{n}x^n}"

"\frac{x^{n+1}}{\frac{x^n\left(n+1\right)}{n}}"

"\frac{x^{n+1}n}{\left(n+1\right)x^n}"

"\frac{nx}{n+1}"

"\lim _{n\to \infty \:}\left(\left|\frac{nx}{n+1}\right|\right)"

"=\left|x\right|\cdot \lim _{n\to \infty \:}\left(\left|\frac{n}{n+1}\right|\right)"

"\left|x\right|\cdot \:1"

"-1<x<1"

"\mathrm{Check\:convergence\:for\:}L=1"

"\lim _{n\to \infty \:}\left(\frac{1}{n}\right)=0"

"\mathrm{Therefore,\:the\:convergence\:interval\:of\:}\sum _{n=1}^{\infty \:}\left(-1^{n-1}\right)x^n\frac{1}{n}\mathrm{\:is\:} -1\le \:x<1"